User Tools

Site Tools


new_measurement_cern

new measurement @ cern

HV scan

Setup

State FE1_24CH_16

Threshold 504

HVscan (HV eff) from 50 → 8000

P=960 mBar T =20 C

run 739145 hangs up (technician working around)

some noise peak (phone calls)

run 739147 started 220 s later stopped because TOP gap tripped Hveff ~ 5380 (Applied 5250)

Severa lost of communication »..?

Run P T HV eff HV app Top Bot
739153 961 20.9 5000 4878 0.10 0.06

DAQ user 2018/02/26 11:27

Measurement with standard Gas

run start end HV Correction I bot Itop Hv eff Eps deps Eps xy Deps xy
739303 9:50 11:52 7400 0.96679 0.63 1.34 7654.195843978593.15 0.93 93.01 0.94
739302 23:25 1:45 7300 0.9671 0.5 1.28 7548.340399131491.66 1.01 91.39 1.03
739304 11:53 13:53 7300 0.96653 0.53 1.07 7552.791946447690.1 2.1 90.12.1
739286 7200 0.9639 0.5 1.31 7469.654528478193.38 0.94 93.04 0.96
739305 16:25 18:35 7200 0.96785 0.15 0.42 7439.16929276231.27 0.42 1.130.4
739306 18:35 20:40 7100 0.96963980.12 0.34 7322.30669574411.55 0.44 1.42 0.43
739301 7000 0.9678 0.4 0.77 7232.899359371887.59 1.23 84.66 1.35
739298 14:55 16:56 6800 0.9655 0.26 0.63 7042.982910409157.25 2.43 53.14 2.45
739299 17:05 19:12 6600 0.9763 0.2 0.45 6760.217146369 12.72 1.66 11.47 1.59
739300 19:13 21:13 6400 0.96663 0.13 0.34 6620.940794306
new_measurement_cern.txt · Last modified: 2021/07/10 23:21 (external edit)